OpenStudy (anonymous):

THE SUM OF THREE CONSECUTIVE ODD INTEGERS IS 195, WHAT ARE THE INTEGERS??? anyone

OpenStudy (anonymous):

your numbers will be represented by x, x + 2, x + 4 x + x + 2 + x + 4 = 195 combine like terms 3x + 6 = 195 3x = 195 - 6 3x = 189 x = 63 x + 2 = 63 + 2 = 65 x + 4 = 63 + 4 = 67 check... 63 + 65 + 67 = 195 195 = 195 (correct) any questions ?